OpenStudy (anonymous):

Which of the following is the quotient of the rational expression 5/x-2 divided by x/x+1?

OpenStudy (phi):

remember dividing by \( \frac{x}{x+1} \) is the same as multiplying by \( \frac{x+1} {x}\)

OpenStudy (anonymous):

@phi then I just get rid of the x/x+1?

OpenStudy (phi):

the first step is write down your problem, using the idea posted up above. what do you get ?

OpenStudy (anonymous):

\[\frac{ 5 }{ x-2}/\frac{ x}{ x+1 }\]

OpenStudy (phi):

that shows division. now change the divide to multiply: flip the 2nd fraction, and multiply \[ \frac{5}{x-2} \cdot \frac{x+1}{x} \] now multiply top times top, and bottom times bottom